OpenStudy (anonymous):

For the given triangles, a = 14.3 m, b = 8.93 m, and α = 50.6 degrees ◦ .Find θ. Answer in units of ◦

OpenStudy (anonymous):



OpenStudy (jdoe0001):

something tells the picture will look a bit more descriptive :)

OpenStudy (jdoe0001):

and since now you know how to take screenshots :)

OpenStudy (anonymous):

I tried you said alt then printscreen right?

OpenStudy (jdoe0001):

no alt needed, just the PrintScr button by itself paste in Paint, save to say .gif or .jpeg

OpenStudy (anonymous):

I can't find the message that you told me

OpenStudy (anonymous):

oh okay will do!

OpenStudy (jdoe0001):

paste to Paint, crop anything you want in the picture, write any extra notes on it, save it and [Attach File]

OpenStudy (anonymous):

my computer doesn't have paint, it was updated and has some artist type of painting but you cant paste anything to it

OpenStudy (jdoe0001):

hmm, ok, any graphic app then

OpenStudy (jdoe0001):

well, you should be able to in the app, start a new image and then do ctrl-v

OpenStudy (anonymous):

alright I think I got it I'll attach in a minute :)

OpenStudy (jdoe0001):

what about this, try this Start > Run > mspaint type that at the Run box

OpenStudy (jdoe0001):

docx? whatever happened to .gif? or jpeg?

OpenStudy (anonymous):

Do you need me to re save it?

OpenStudy (jdoe0001):

:)

OpenStudy (anonymous):

oh okay thanks, now what can I do to find the answer?

OpenStudy (jdoe0001):

notice the triangles, they're similar, that means their angles and sides have a ratio relationship

OpenStudy (jdoe0001):

on the bigger one, there are 2 angles, since they're both right-triangles one angle is 90 degrees, the other is 50.6 degrees so the missing internal angle will be?

OpenStudy (anonymous):

39.4

OpenStudy (anonymous):

degrees

OpenStudy (jdoe0001):

yes, thus using ratio relationship, we can say that \(\bf \cfrac{39.4^o}{\theta}=\cfrac{14.3}{8.93}\) then just solve for \(\bf \theta\)

OpenStudy (anonymous):

63.0929451288?

OpenStudy (anonymous):

degrees

OpenStudy (jdoe0001):

hmm, if you look at the picture, 63 degrees will make a bit wider than that so is smaller than that

OpenStudy (jdoe0001):

\(\bf \cfrac{39.4^o}{\theta}=\cfrac{14.3}{8.93} \implies \cfrac{39.4\times 8.93 }{14.3}=\theta\)

OpenStudy (anonymous):

24.6043356643

OpenStudy (jdoe0001):

:)

OpenStudy (anonymous):

Great thanks sooo much !!! I have one more question on my hw and someone else tried to help me but I remained confused. I really just need the equation on how to solve it.:)

OpenStudy (jdoe0001):

ok

OpenStudy (anonymous):

Hey the website for my hw says that 24.6043356643 is incorrect

OpenStudy (jdoe0001):

----> Answer in units of ◦ <---- so I gather it'd be a rounded up answer, of 25 or 24.6

OpenStudy (jdoe0001):

you know is easier to use .jpeg or .gif over .docx :) but anyhow

OpenStudy (anonymous):

oh sorry

OpenStudy (jdoe0001):

this one?

OpenStudy (anonymous):

yes

OpenStudy (jdoe0001):

well, I see the rhombus I see the angle so, what's sought after?

OpenStudy (anonymous):

Find the distance from A to B. Answer in units of cm

OpenStudy (anonymous):

So what do I do next?